OpenStudy (anonymous):

Find all zeroes. f(x)=(2x-1)(x-5) PLEASE HELP ME STEP BY STEP ON HOW TO DO IT.*

OpenStudy (anonymous):

(2x-1)(x-5) will be equal to zero if either (2x-1)=0 or (x-5)=0, so: 2x-1 = 0 x = 1/2 Or x-5 = 0 x = 5

OpenStudy (anonymous):

ok @M4thM1nd

OpenStudy (anonymous):

how is x=1/2 how did you find that

OpenStudy (anonymous):

2x-1 = 0 2x = 1 x = 1/2

OpenStudy (anonymous):

but would'ntt it be 2x-1=1

OpenStudy (anonymous):

No, first you have: 2x - 1 = 0, now sum 1 in both sides and we get: 2x -1 + 1 = 0 + 1 2x = 1, now divide both sides by 2, we get: 2x/2 = 1/2 x = 1/2

OpenStudy (anonymous):

i dont understand but thank you i know im annoying